OpenStudy (anonymous):

Solve the quadratic equation. (x - 4)2 = 36

OpenStudy (amistre64):

x = 10; x = -2

OpenStudy (anonymous):

thanks. i seriously need you as a tutor ! lol

OpenStudy (amistre64):

:) this aint tutor; this is more human calculator lol

OpenStudy (anonymous):

(x-4)^2 ??????????

OpenStudy (amistre64):

6^2 = 36 (-6)^2 = 36 10-4 = 6 -2-4 = -6